Suppose the tube in the pervious problem is kept vertical with A upward but the other conditions remain the same. The separation between the cross sections at A and B is 15/16 cm. Repeat parts (a), (b) and (c) of the previous problem. Take g = 10 m s-2.

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol. From Bernoulli’s equation, 1/2 4ρvA2 + pghA + pA = 1/2 pvB2 + pghB + pB. ⇒PA – PB = (1/2) p (vB2 – vA2) + pg (hB – hA)